# Symmetry-aware Reinforcement Learning for Robotic Assembly under Partial Observability with a Soft Wrist

### v1

@@ I INTRODUCTION
A peg-in-hole task often abstracts an insertion operation in robotic assembly [1]. This task is widely explored [1] but is still challenging due to the contact-richness and tight clearance between parts. Most approaches involve rigid robots with force control [2], sending high-frequency control signals for safe operation. As an alternative, similar to recent studies [3, 4, 5], this study utilizes a physically soft robot for contact-rich tasks, given its softness can tolerate contacts and low-frequency control signals without sacrificing safety.
@@
  • We conduct experimental evaluations across five symmetric peg shapes in simulation using MuJoCo [8]. The results show that our proposed agent can be positively comparable to a state-based agent. Moreover, the learned policy can be generalized from one shape to another. In particular, the sample efficiency allows us to learn the task directly on a real robot using 100 demonstrations within 3 hours.
@@ II-A Pose Estimation in Soft Robots
Motion capture systems are often used to measure soft robots’ poses [6, 9, 10]. Soft robotic assembly scenarios [11, 12] also assumed a full observability setting in which the peg-to-hole pose is measured using an external motion capture setup. As alternative approaches, embedded bending sensors [13], air-pressure [14] or, multi-modal sensors [15, 16] have been used. Instead, this study uses F/T sensors to indirectly obtain the contact states of the robot. A relevant study uses an F/T sensor to estimate an elastic rod’s shape with a model-based approach [17]. In contrast, we do not require a model assumption and instead employ a model-free deep reinforcement learning (DRL) approach considering the POMDP setting.
@@ II-C Symmetry-aware Policy Learning
Exploiting the domain symmetry can benefit robot learning, especially when samples are often expensive to collect with real robot systems. Under full observability, recent studies such as [21, 22, 23, 24] introduced various symmetry-aware RL agents that encode the domain symmetry directly into their network architectures. In particular, [25] instead uses the symmetry to transform the state and feed both the original and transformed states to the actor and the critic, making an actor-critic agent symmetry-aware. In contrast to the fully observable setting, incorporating symmetry is under-explored under partial observability, which is considered more difficult to solve. A recent work in this direction is [26], which enforces symmetry by directly constructing every component of an actor-critic agent with equivariant (symmetry-aware) neural networks [27]. In contrast, we use no specialized neural networks, and instead indirectly enforce the symmetry through auxiliary losses and additionally use the domain symmetry for data augmentation, applied to a recurrent Soft Actor-Critic (SAC) [28] agent. Moreover, the way we use auxiliary losses to enforce the symmetry differs from the common practices in the POMDP literature, in which auxiliary tasks are often used to learn better history representations through reconstruction tasks [29] or prediction tasks [30].
@@ III-B Notations
We use $G$ to denote a set of group elements $g$; each defines a transformation. A special element of $G$ is the identity transformation $e$, which leaves the input unchanged. This work focuses on 2D transformations; therefore, $G$ refers to planar reflections and rotations. For the planar reflections, we use $F^{x}=\{e,r^{x}\}$ for the X-axis reflection, with $r^{x}$ defining the reflection. We denote the result of applying $G$ on a 2D point $p_{0}=(x,y)$ as $G*p_{0}$, which results in a new set, whose elements are $g*p_{0}$ for each $g\in G$. For instance, $F^{x}*p_{0}=\{e*p_{0},r^{x}*p_{0}\}=\{p_{0},p_{1}\}$ (see Fig. 2). Similarly, $F^{y}=\{e,r^{y}\}$ refers to the Y-axis reflection. For convenience, we use $F^{xy}$ to shorthand for $F^{x}*F^{y}$, which combines the two transformations sequentially, i.e., $F^{xy}*p_{0}=\{p_{0},p_{1},p_{4},p_{5}\}$. For rotational transformations, we use $R^{n}$ to refer to a set of $n$ planar rotations $\{0,2\pi/n,4\pi/n,\dots,2\pi(n-1)/n\}$ ($e$ corresponds to a rotation of 00 rad). Specifically, a transformation element $g\in R^{n}$ will rotate $p_{0}$ counter-clockwise around the origin. For instance, $R^{4}*p_{0}=\{p_{0},p_{2},p_{4},p_{6}\}$. Similar to $F^{xy}$, we can combine $F^{xy}$ and $R^{4}$, i.e., $F^{xy}*R^{4}*p_{0}=(p_{0},\dots,p_{7})$. Finally, we use $|G|$ to denote the group size of $G$, e.g., $|R^{n}|=n$, $|F^{x}|=|F^{y}|=2$, $|F^{xy}|=4$, and $|F^{xy}*R^{4}|=8$.
@@ VII-A Setup
Simulation results allow us to perform learning on hardware, using a soft wrist [39] mounted on a UR5e robot (see Fig. 9) and a round and a square peg, each with a 1-mm clearance. The integrated F/T sensor (parameters in Fig. 9) is located before the soft wrist, which measures the F/T signals applied to the entire arm. This differs from the simulation setting, in which only the F/T signals applied to the peg are measured. The peg’s initial position is randomized to touch the hole’s surface to prevent mid-air vibration when moving. We filter the signals with a low-pass filter (50-Hz cut-off frequency and 500-Hz sampling frequency). We also remove the F/T offsets when starting an episode due to sensor drift.
@@ IX CONCLUSIONS
This study enhances a recurrent SAC agent for solving symmetric POMDPs by integrating data augmentation and auxiliary losses for training an actor-critic DRL agent. Our method’s versatility extends beyond the peg-in-hole task with a soft wrist and can be adapted for other scenarios. A clear limitation of our method is the requirement of near-perfect symmetries for optimal performance, yet minor imperfections in real-robot experiments, like surface unevenness or inconsistency, did not significantly affect the performance. Also, recent studies [40, 41] demonstrate the surprising benefits of leveraging symmetries in imperfect settings, which can potentially apply to our POMDP context.

### v2

@@ I INTRODUCTION
A peg-in-hole task often abstracts an insertion operation in robotic assembly [1]. This task is widely explored [1] but is still challenging due to the contact-richness and tight clearance between parts. Most approaches involve rigid robots with force control [2], sending high-frequency control signals for safe operation. As an alternative, physically soft robots can be used for contact-rich tasks, given their softness can tolerate contacts and low-frequency control signals without sacrificing safety [3, 4, 5, 6].
@@
  • We conduct experimental evaluations across five symmetric peg shapes in simulation using MuJoCo [9]. The results show that our proposed agent can even be positively comparable to a state-based agent. Moreover, the learned policy can be generalized from one shape to another. In particular, the sample efficiency allows us to learn the task directly on a real robot using 100 demonstrations within 3 hours.
@@ II-A Pose Estimation in Soft Robots
Motion capture systems are often used to measure soft robots’ poses [7, 10, 11]. Soft robotic assembly scenarios [12, 13] also assumed a full observability setting in which the peg-to-hole pose is measured using an external motion capture setup. As alternative approaches, embedded bending sensors [14], air-pressure [15] or, multi-modal sensors [16, 17] have been used. Instead, this study uses F/T sensors to obtain the contact states of the robot indirectly. A relevant study uses an F/T sensor to estimate an elastic rod’s shape with a model-based approach [18]. In contrast, we do not require a model assumption and instead employ a model-free deep reinforcement learning (DRL) approach considering the POMDP setting.
@@ II-C Symmetry-aware Policy Learning
Exploiting the domain symmetry can benefit robot learning, especially when samples are often expensive to collect with real robot systems. Under full observability, recent studies such as [22, 23, 24] introduced various symmetry-aware RL agents that encode the domain symmetry directly into their network architectures. In particular, [25] instead uses the symmetry to transform the state and feed both the original and transformed states to the actor and the critic, making an actor-critic agent symmetry-aware. In contrast to the fully observable setting, incorporating symmetry is under-explored under partial observability, which is considered more difficult to solve. A recent work in this direction is [26], which enforces symmetry by directly constructing every component of an actor-critic agent with equivariant (symmetry-aware) neural networks [27]. In contrast, we use no specialized neural networks, and instead indirectly enforce the symmetry through auxiliary losses and additionally use the domain symmetry for data augmentation, applied to a recurrent Soft Actor-Critic (SAC) [28] agent. Moreover, the way we use auxiliary losses to enforce the symmetry differs from the common practices in the POMDP literature, in which auxiliary tasks are often used to learn better history representations through reconstruction tasks [29] or prediction tasks [30].
@@ III-B Notations
We use $G$ to denote a set of group elements $g$; each defines a transformation. A special element of $G$ is the identity transformation $e$, which leaves the input unchanged. This work focuses on 2D transformations; therefore, $G$ is limited to planar reflections and rotations. For the planar reflections, we use $F^{x}=\{e,r^{x}\}$ for the X-axis reflection, with $r^{x}$ defining the reflection. We denote the result of applying $G$ on a 2D point $p_{0}=(x,y)$ as $G*p_{0}$, which results in a new set, whose elements are $g*p_{0}$ for each $g\in G$. For instance, $F^{x}*p_{0}=\{e*p_{0},r^{x}*p_{0}\}=\{p_{0},p_{1}\}$ (see Fig. 2). Similarly, $F^{y}=\{e,r^{y}\}$ refers to the Y-axis reflection. For convenience, we use $F^{xy}$ to shorthand for $F^{x}*F^{y}$, which combines the two transformations sequentially, i.e., $F^{xy}*p_{0}=\{p_{0},p_{1},p_{4},p_{5}\}$. For rotational transformations, we use $R^{n}$ to refer to a set of $n$ planar rotations $\{0,2\pi/n,4\pi/n,\dots,2\pi(n-1)/n\}$ ($e$ corresponds to a rotation of 00 rad). Specifically, a transformation element $g\in R^{n}$ will rotate $p_{0}$ counter-clockwise around the origin. For instance, $R^{4}*p_{0}=\{p_{0},p_{2},p_{4},p_{6}\}$. Similar to $F^{xy}$, we can combine $F^{xy}$ and $R^{4}$, i.e., $F^{xy}*R^{4}*p_{0}=(p_{0},\dots,p_{7})$. Finally, we use $|G|$ to denote the group size of $G$, e.g., $|R^{n}|=n$, $|F^{x}|=|F^{y}|=2$, $|F^{xy}|=4$, and $|F^{xy}*R^{4}|=8$.
@@ VII-A Setup
We perform learning on hardware, using a soft wrist [39] mounted on a UR5e robot (see Fig. 9) and a round and a square peg, each with a 1-mm clearance. The F/T sensor installed in UR5e (parameters in Fig. 9) is located before the soft wrist, which measures the F/T signals applied to the entire arm. This differs from the simulation setting, in which only the F/T signals applied to the peg are measured. The peg’s initial position is randomized to touch the hole’s surface to prevent mid-air vibration when moving. We filter the signals with a low-pass filter (50-Hz cut-off frequency and 500-Hz sampling frequency). We also remove the F/T offsets when starting an episode due to sensor drift.
@@ IX CONCLUSIONS
This study enhances a recurrent SAC agent for solving symmetric POMDPs by integrating data augmentation and auxiliary losses. Our method’s versatility extends beyond the peg-in-hole task and can be adapted for other scenarios. A limitation of our method is the requirement of near-perfect symmetries for optimal performance. Yet, minor imperfections in real-robot experiments, like surface unevenness or inconsistency, did not significantly affect the performance. Also, a recent study shows the benefits of leveraging symmetries in imperfect settings [40] and the improved robustness informing of the contact states using tactile sensors [41], which can potentially apply to our POMDP context.